# A Simple DropConnect Approach to Transfer-based Targeted Attack

### v1

@@ 6 Concluding Remarks
In this paper, we first show the perturbation co-adaptation issue in existing transfer-based attack methods, and then propose MCD that creates diverse variants of surrogate model at each iteration to improve transferability. MCD brings in more diversification across the variants while maintaining the utility of each variant. Extensive experiments demonstrate the superior performance of MCD.
@@
Appendix A Visualization of Adversarial Examples
@@
Figs. 6 and 7 visualize the attention heatmap of the adversarial examples, with surrogate model ResNet-50 and target model ResNet-18. We observe that the attention heatmaps w.r.t. MCD on the surrogate model are more similar to the heatmaps on target model, compared with other attack methods. This in part explains that the generated adversarial example of MCD is more transferable to the target model.
@@
Appendix B Additional Experimental Results
B.1 Additional Results in Pilot Study
We present more pilot study results discussed in Sec. 4.1.
(a) DenseNet-121 $\to$ ResNet-18
(b) DenseNet-121 $\to$ MobileNet-v2
Figure 8: ASRs(%) with increasing mask ratio.
B.2 Initial Results of Applying Dropout to Perturbation and Intermediate Features
Tab. 5 reports the results of applying random masking to perturbation as well as to intermediate features. We see that the Dropout idea generally helps with increased ASR, but does not achieve the best performance. Notice that CFM is also implemented in combination with RDI, which we write it explicitly as RDI+CFM in the table.
We also include the empirical result of Ghost Networks (GN) [22]. It perturbs only skip connections for residual networks and also uses one new model at each iteration. For a fair comparison, we combine it with RDI and pick a larger iteration number for optimization. However, its performance is still much lower. Compared with RDI+DropF, we can see that the additional randomness of applying Dropout to each layer is beneficial to transfer-based targeted attack.
Method DI GN DI+DropP DI+DropF RDI RDI+DropP RDI+DropF RDI+CFM
ASR(%) 26.8 29.7 38.7 60.0 42.0 48.7 63.0 66.1
Table 5: Average ASRs of applying Dropout onto perturbation and intermediate features, denoted as +DropP and +DropF, respectively. Results are the average over 16 target models on the ImageNet-compatible dataset, with ResNet-50 as surrogate model. Detailed experimental setting can be found in Section 5.1.
B.3 Ablation Study on DropConnect Probability
Experimental results of varying DropConnect probabilities are shown in Fig. 9, indicating a stable performance of MCD across a range of different parameter values.
Figure 9: ASRs(%) averaged over 16 different models with varying DropConnect probabilities, using ResNet-50 as the surrogate model.
B.4 Additional Attack Results
Tab. 6 and Tab. 7 present additional attack results. The setting is the same as described in Sec. 5.1. Our method MCD outperforms other baselines by a large margin. Furthermore, in the setting where CLIP is used as a surrogate model, MCD again results in a notable improvement. Yet, the ASRs are still relatively low, indicating that there is much room for further enhancement, particularly in the case of transferring from CLIP to CNN-based models.

### v2

@@ 5 Concluding Remarks
In this paper, we reveal the over-reliance issue in existing transfer-based attacks, where adversarial examples depend excessively on a small subset of model parameters. To alleviate this, we propose RaPA, which randomly prunes surrogate parameters during optimization. We show that the expected effect of random pruning equals adding an importance-equalization regularizer, thereby reducing parameter over-reliance and improving transferability. Extensive experiments on CNN and Transformer architectures confirm the effectiveness and stability of RaPA.
@@
Appendix A More Related Work
The first class applies transformations to the input images, thereby diversifying the input patterns [52, 57, 9, 1, 48, 46, 2]. Diverse Inputs (DI) [52] is one of the representative methods. At each iteration, it applies random and differentiable transformations (e.g., random resizing) to the input image and then maximizes the loss function w.r.t. the transformed inputs. Resized Diverse Inputs (RDI) [57] extends DI by resizing the transformed image back to its original size. The Translation-Invariant (TI) attack method [9] optimizes perturbations across a set of translated images, making the adversarial example less sensitive to the surrogate model. Block Shuffle and Rotate (BSR) [46] divides the input image into blocks, followed by shuffling or rotation, and Object-based Diversity Inputs (ODI) [1] generates an adversarial example on a 3D object and leads the rendered image to being classified as the target class. Admix [48] mixes the input image with random samples from other classes. Clean Feature Mixup (CFM) [2] extends Admix to high-level feature space and introduces two types of competing noises to guide adversarial perturbations. Building upon it, Feature Tuning Mixup (FTM) [26] further introduces learnable and attack-specific feature perturbations that combine random and optimized noises in the feature space, achieving SoTA performance in transfer-based targeted attacks.
Another line of work focuses on improving the surrogate model itself through additional training to enhance the adversarial transferability. Dark Surrogate Model (DSM) [53] trains a surrogate model using dark knowledge distilled from a teacher model and further augments the training data with mixing augmentations such as CutMix and Mixup, thereby enriching soft supervision and improving transferability. SASD-WS [50] introduces Sharpness-Aware Self-Distillation (SASD) to flatten the loss landscape and combines it with Weight Scaling (WS) to approximate model ensembling. While these approaches require retraining surrogate models, our RaPA works in a training-free manner, yet achieves comparable or even superior transferability, and can be seamlessly integrated with such training-based frameworks.
Closely related to the present work is self-ensemble [19, 54, 25, 33], which creates multiple models from only one surrogate model. Transfer-based self-ensemble (T-sea) [19] locally apply various transformations onto the input image to improve such diversity while preserving the structure of imagetargets the task of object detection, and the self-ensemble method in [33] specifically considers vision Transformer as surrogate model and is denoted as SE-ViT in this paper. Ghost Network [25] perturbs surrogate model to create a set of new models and then samples one model from the set at each iteration. Masking Unimportant Parameters (MUP) [54] drops out unimportant parameters according to a predefined Taylor expansion-based metric.
Appendix B Estimation and Approximation
B.1 Parameter Importance Estimation and Pruning by Importance
Computing the second-order derivative defined in Eq. (3) is generally costly. A more compact approximation is obtained via first-order expansion [32], which reduces to
$\mathcal{I}(\theta_{i})=\frac{\partial^{2}\mathcal{L}f(x_{\mathrm{adv}})}{\partial\theta_{i}^{2}}\theta_{i}^{2}\approx\left(\frac{\partial\mathcal{L}f(x_{\mathrm{adv}})}{\partial\theta_{i}}\theta_{i}\right)^{2}.$ (7)
In our pilot study, we adopt the method in DepGraph [11] to perform the pruning, which explicitly models inter-layer dependencies and comprehensively groups coupled parameters for pruning.
B.2 Derivation of the Masked Loss Approximation in Eq. (4)
Let $\theta$} represent the entire set of model parameters. We define a random binary mask $\mathcal{M}\in\{0,1\}^{|\theta|}$, where each entry is independently sampled from a Bernoulli distribution: $\mathcal{M}_{i}\sim\text{Bernoulli}(1-p)$. A second-order Taylor expansion gives
$\mathcal{L}\bigl(f(x_{\mathrm{adv}};\mathcal{M}\odot\theta)\bigr)\approx$ $~\mathcal{L}\bigl(f(x_{\mathrm{adv}};\theta)\bigr)+\sum_{i}g_{i}\Delta_{i}+\tfrac{1}{2}\sum_{i,j}H_{ij}\Delta_{i}\Delta_{j},$
where $\Delta=(\mathcal{M}-\mathbf{1})\odot\theta$, $g_{i}=\partial\mathcal{L}\bigl(f(x_{\mathrm{adv}};\theta)\bigr)/\partial\theta_{i}$ and $H_{ij}=\partial^{2}\mathcal{L}\bigl(f(x_{\mathrm{adv}};\theta)\bigr)/(\partial\theta_{i}\partial\theta_{j})$ evaluated at $\theta$. Since $p$ is taken to be small and close to zero, we have $\mathbb{E}[\Delta_{i}]\approx 0$. And the mask entries are independent, so we further have $\mathbb{E}[\Delta_{i}\Delta_{j}]\approx 0$ for $i\neq j$ while $\mathbb{E}[\Delta_{i}^{2}]=p(1-p)\theta_{i}^{2}$. Taking expectations yields
$\mathbb{E}_{\mathcal{M}}\!\left[\mathcal{L}\bigl(f(x_{\mathrm{adv}};\mathcal{M}\odot\theta)\bigr)\right]\approx\mathcal{L}\bigl(f(x_{\mathrm{adv}};\theta)\bigr)+\frac{p(1-p)}{2}\sum_{i}H_{ii}\,\theta_{i}^{2},$ (8)
which is the expression in Eq. (4).
B.3 Detailed Gini Coefficients Computation Procedure
We compute the Gini coefficients based on the parameter importance values $\mathcal{I}(\theta_{i})$ defined in Eq. (3). For each layer $l$ in the surrogate model, we first collect all parameter importance values $\{\mathcal{I}^{l}(\theta_{i})\}_{i=1}^{n}$. To better highlight the disparity among parameters, we further apply an exponential scaling to these importance values before computing the Gini coefficient:
$\tilde{\mathcal{I}}^{l}(\theta_{i})=\exp\left(\mathcal{I}^{l}(\theta_{i})\right),$ (9)
where $\tilde{\mathcal{I}}^{l}(\theta_{i})$ denotes the scaled importance. This transformation magnifies the relative differences among parameters, enabling a clearer assessment of importance inequality.
Given the scaled importance values $\{\tilde{\mathcal{I}}^{l}(\theta_{1}),\tilde{\mathcal{I}}^{l}(\theta_{2}),\dots,\tilde{\mathcal{I}}^{l}(\theta_{n})\}$ in layer $l$, we compute the Gini coefficient as:
$\text{Gini}^{l}=\frac{2\sum_{i=1}^{n}i\tilde{v}_{(i)}}{n\sum_{i=1}^{n}\tilde{v}_{(i)}}-\frac{n+1}{n},$ (10)
where $\tilde{v}_{(i)}$ represents the $i$-th smallest value after sorting in ascending order. A smaller Gini coefficient indicates a more balanced distribution of parameter importance within the layer.
Finally, we aggregate the Gini coefficients across layers to obtain both the type-wise and overall average results:
$\text{Gini}_{\text{avg}}=\frac{1}{|\mathbb{L}|}\sum_{l\in\mathbb{L}}\text{Gini}^{l},$ (11)
where $\mathbb{L}$ denotes the set of all layers containing learnable parameters. This process allows us to quantitatively evaluate the degree of importance imbalance and verify that the proposed random parameter pruning strategy effectively reduces the model’s over-reliance on a small subset of dominant parameters.
Appendix C Visualization of Adversarial Examples
@@
Figs. 4 and 5 visualize the attention heatmap of the adversarial examples, with surrogate model ResNet-50 and target model ResNet-18. We observe that the attention heatmaps w.r.t. RaPA on the surrogate model are more similar to the heatmaps on target model, compared with other attack methods. This in part explains that the generated adversarial example of MCD is more transferable to the target model.
@@
Appendix D Additional Experimental Results
D.1 Result of masking the most important parameters
Method DI GN RDI MMP(1%) MMP(0.1%) MMP(0.01%) MMP(0.001%)
ASR(%) 26.8 29.7 42.0 0.2 27.7 36.7 38.9
Table 8: Average ASRs when masking the most important parameters (MMP) in the surrogate model at mask ratios of 1 %, 0.1 %, 0.01 %, and 0.001 %. Results are the average over 16 target models on the ImageNet-compatible dataset, with ResNet-50 as surrogate model. Detailed experimental setting can be found in Section 4.1.
Using the approximation in Eq. 7, we mask the most important parameters. The results are reported in Table 8. As analyzed in Section 3, masking the most important parameters degrades the model’s attacking capability. A high masking ratio (e.g., 1 %) causes the attack to fail, whereas a low ratio (e.g., 0.001 %) leaves transferability unaffected, so the ASRs converge to the baseline RDI.
We also include the empirical result of Ghost Networks (GN) [25]. It perturbs only skip connections for residual networks and also uses one new model at each iteration. For a fair comparison, we combine it with RDI and pick a larger iteration number for optimization. However, its performance is still much lower.
D.2 Ablation Study on DropConnect Probability
Experimental results of varying DropConnect probabilities are shown in Fig. 6, indicating a stable performance of RaPA across a range of different parameter values.
Figure 6: ASRs(%) averaged over 16 models with varying DropConnect probabilities, using ResNet-50 as surrogate .
D.3 Additional Attack Results
Tab. 9 and Tab. 10 present additional attack results. The setting is the same as described in Sec. 4.1. Our method RaPA outperforms other baselines by a large margin. Furthermore, in the setting where CLIP is used as a surrogate model, RaPA again results in a notable improvement. Yet, the ASRs are still relatively low, indicating that there is much room for further enhancement, particularly in the case of transferring from CLIP to CNN-based models.
@@ D.3 Additional Attack Results
D.4 Model Diversity and Utility
(a) Diversity
(b) Utility
Figure 7: Diversity and Utility with increasing mask ratio.
We show the diversity and utility of constructed models. Here diversity is measured by first computing the mean KL-divergence between the output distributions of each pair of model variants for an input and then averaging these values over the ImageNet-compatible dataset. Similarly, the utility is quantified by the mean top-1 accuracy over the model variants and the dataset. We use the same surrogate model ResNet-50 as in the ablation study.
The results are presented in Figure 7. As DropConnect probability $p$ increases, the diversity among different variants improves while the utility decreases. As shown in Figure 6 in the supplementary material, the average ASR using the same surrogate model increases with $p\in[0.01,0.05]$ and then decreases with $p\in[0.05,0.09]$, while the highest ASR is obtained at $p=0.05$. This verifies our discussion in Section 3.3 that the attack performance is affected by both diversity and utility. We conclude that the better performance of RaPA is due to the improved diversification across the variants while keeping each variant sufficiently useful.
Ethical Statement
While our work is intended for research and defense purposes, it could be misused by malicious actors to craft black-box adversarial attacks against safety-critical systems such as autonomous-driving or medical models. By demonstrating that these systems remain vulnerable even when attackers have no knowledge of their internal details, we aim to alert service providers and researchers to this threat and to accelerate the development of more robust deep-learning defenses.